# Concurrent-Learning Based Relative Localization in Shape Formation of Robot Swarms

### v1

@@ IV-A Relative Localization Problem
For each robot $i\in\mathcal{V}$ and $j\in\mathcal{N}_{i}$, denote the relative position between robots $i$ and $j$ at time $t_{k}=k\Delta t$ as
$p_{ij}(t_{k})=p_{i}(t_{k})-p_{j}(t_{k}),$
where $k\in\mathbb{N}$ is the number of samples, and $\Delta t$ is the sample interval. Here, $p_{ij}(t_{k})$ is related to its initial relative position $p_{ij}(t_{0})$ satisfying
$p_{ij}(t_{k})=p_{ij}(t_{0})+z_{ij}(t_{k}),$ (1)
where $z_{ij}(t_{k})$ is defined as $z_{ij}(t_{k})=z_{i}(t_{k})-z_{j}(t_{k})$, $z_{i}(t_{k})$ is the displacement measured by IO of robot $i$ from $t_{0}=0$ to $t_{k}$. Note that $z_{ij}(t_{k})$ is obtainable through the wireless communication between robots $i$ and $j$.
Next, we can formulate the relative localization problem, i.e., the determination of $p_{ij}(t_{k})$, as an estimation problem of parameter $p_{ij}(t_{0})$. Accordingly, we can define
$y_{ij}(t_{k})=u_{ij}^{T}(t_{k})p_{ij}(t_{0}),$ (2)
@@
$y_{ij}(t_{k})=u_{ij}^{T}(t_{k})p_{ij}(t_{k})-u_{ij}^{T}(t_{k})z_{ij}(t_{k}).$ (3)
For the first term in (3), it follows from the cosine theorem that
$u_{ij}^{T}(t_{k})p_{ij}(t_{k})=\frac{1}{2}\left(d_{ij}^{2}(t_{k+1})-d_{ij}^{2}(t_{k})-u_{ij}^{2}(t_{k})\right),$ (4)
where $d_{ij}$ represents the relative distance between robots $i$ and $j$. The value of $d_{ij}$ in (4) can be measured by UWB. Thus, the first term $u_{ij}^{T}p_{ij}$ in (3) is also obtainable. For better understanding of (4), an illustrative example is introduced in Fig. 3. As shown in Fig. 3A, it can be seen that the relative displacement $u_{ij}(t_{k})$ and the relative positions between the two robots at $t_{k}$ and $t_{k+1}$ form a triangle. For the second term in (3), since both $u_{ij}$ and $z_{ij}$ are directly obtainable, the second term $u_{ij}^{T}z_{ij}$ in (3) is also obtainable. With the fact that both $y_{ij}(t_{k})$ and $u_{ij}(t_{k})$ can be measured, adaptive parameter update law can be designed to estimate the initial relative position $p_{ij}(t_{0})$. As a result, the relative localization task can be achieved.
Denote $\hat{p}_{ij,0}(t_{k})$ as an estimation of the initial relative position $p_{ij}(t_{0})$ and $\tilde{p}_{ij,0}(t_{k})=\hat{p}_{ij,0}(t_{k})-p_{ij}(t_{0})$ as the estimation error. Accordingly, we can define the innovation $\epsilon_{ij}(t_{k})$ as
@@
The innovation $\epsilon_{ij}(t_{k})$ will be used later to design the concurrent-learning based estimator in Section IV-C.
Based on the estimation problem modeled as (2), methods proposed in [12], [17] and [38] attempt to estimate the initial relative position $p_{ij}(t_{0})$. However, all these existing results require that relative displacement $u_{ij}(t)$ satisfies the PE condition. To satisfy this condition, the persistent noise or periodic signal are adopted in the control command of the robot, which are not conducive to the expansion of massive robot swarms.
@@ IV-B Data Collection Strategy
Figure 3: Geometric relationship between the displacements and distance measurements of the two robots. A. one sample interval $\Delta t$. B. $h$ sample interval $\Delta t$
In the relative localization problem, $u_{ij}(t_{k})$ and $d_{ij}(t_{k})$ are two critical values to determine $p_{ij}(t_{0})$. If the sample interval $\Delta t$ is too small, the measurements of $u_{ij}(t_{k})$ may be drowned in noise in practice. At this time, it is necessary to use tandem measurements to address this issue. To achieve this, we propose the following data collection strategy.
Whenever two robots start localization, a data collection takes place between them. Let $t_{{\rm c},m}$ denote the $m$-th ($m=1,\ldots,\varsigma$) data collection time. The data collection process begins at $t_{{\rm c},1}$ and terminates at $t_{{\rm c},\varsigma}$. The interval between the two consecutive collections is $h\Delta t,h\in\mathbb{N}$, i.e., $t_{{\rm c},m+1}-t_{{\rm c},m}=h\Delta t$, as shown in Fig. 3. The heuristic values of $h$ and $\varsigma$ will be discussed later in Section IV-D. In this way, the innovation $\epsilon_{ij}(t_{k})$ in (5) can be rewritten as
$\epsilon_{ij}(t_{c,m})=u_{ij}^{T}(t_{c,m})\hat{p}_{ij,0}(t_{k})-y_{ij}(t_{c,m}),$ (6)
with
$y_{ij}(t_{c,m})=$ $\frac{1}{2}\left(d_{ij}^{2}(t_{c,m+1})-d_{ij}^{2}(t_{c,m})-u_{ij}^{2}(t_{c,m})\right)$
$-u_{ij}^{T}(t_{c,m})z_{ij}(t_{c,m}),$
where $z_{ij}(t_{{\rm c},m})=\sum_{\iota=1}^{m}u_{ij}(t_{{\rm c},\iota})$. The innovation $\epsilon_{ij}(t_{c,m})$ will be used later in Section IV-C. Note that the tandem data collection strategy is helpful for increasing the modulus of the relative displacement $u_{ij}(t_{c,m})$, which is beneficial for reducing relative error in noisy measurements. This trick helps to improve estimation accuracy in real application.
@@ IV-C Concurrent-Learning based Estimator
Now, we can design the concurrent-learning based estimator to update $\hat{p}_{ij,0}(t_{k})$ for the initial relative position $p_{ij}(t_{0})$ as
@@
$\eta=\frac{\lambda_{\min}(S_{ij})}{(\lambda_{\max}(U_{ij}(t_{k}))+\lambda_{\max}(S_{ij}))^{2}},$ (8)
where matrices $U_{ij}(t_{k})$ and $S_{ij}$ are defined as $U_{ij}(t_{k})=u_{ij}(t_{k})u_{ij}^{T}(t_{k})\in\mathbb{R}^{2\times 2}$ and $S_{ij}=R_{ij}R_{ij}^{T}\in\mathbb{R}^{2\times 2}$, respectively. Here, $R_{ij}=[u_{ij}(t_{c,1}),...,u_{ij}(t_{c,\varsigma})]$. Note that (7) represents the updates at time instant $t_{k+1}$ and employs both the present measurements $d_{ij}(t_{k+1})$, $d_{ij}(t_{k})$, $u_{ij}(t_{k})$ and the historical measurements $d_{ij}(t_{c,m})$, $d_{ij}(t_{c,m+1})$, $u_{ij}(t_{c,m})$ in its calculations. All these measurements are available at $t_{k+1}$.
@@
Recall that $z_{ij}(t)$ is the relative displacement which is available through the IO measurement and inter-robot communication. In the next theorem, we will show that the convergence of the proposed estimator (7) can be guaranteed if the recorded data matrix $S_{ij}$ satisfies the following assumption.
@@ IV-D Localization Enhancement Motion
In real applications, nevertheless, we still need to address the questions on how to design the relative motion trajectories of two robots and how to collect the measurement data, i.e. how to choose $\varsigma$ and $h$ such that the convergence speed of estimator (7) is the largest.
According to Theorem 1, to improve the convergence speed of the estimator, the ratio of the minimum and maximum eigenvalue $\frac{\lambda_{\min}(S_{ij})}{\lambda_{\max}(S_{ij})}$ of the recorded data matrix $S_{ij}$ should be as large as desired. To achieve this, we design the circular motion velocity of each robots $i$ as
$v_{i}$ $=[r_{i}\mathrm{cos}(w_{i}t),r_{i}\mathrm{sin}(w_{i}t)],$ (11)
where $w_{i}=1/i$ is the angular velocity. For the two robot $i$ and $j$, select the measurement interval time $h\Delta t$ and the data set number $\varsigma$ which satisfy
$\mathrm{mod}(2\pi,h\Delta t)$ $=0,$ (12a)
$\varsigma$ $=ij\varsigma_{0},$ (12b)
@@
Theorem 2 shows that with careful design of the relative motion of the two robots, the proposed estimator has the largest convergence rate. In the practical applications, the ratio $\frac{\lambda_{\min}(S_{ij})}{\lambda_{\max}(S_{ij})}$ can be calculated at each data measurement time instant $t_{c,m}$. When $\frac{\lambda_{\min}(S_{ij})}{\lambda_{\max}(S_{ij})}$ is greater than a threshold $\lambda_{0}$, the data collection is sufficient, there is no need to maintain the relative motion between the two robots and the estimator will converge exponentially. This result directly supports our design of localization enhance command in Section VI B.
@@ V Behavior-Based Shape Formation
Previously, we have shown how to achieve relative localization with only distance and displacement measurements. In this section, we will address the problem how to achieve shape formation control through behavior based control strategy. To achieve this, we first define and parameterize the desired formation shape. Then, a consensus based shape localization agreement protocol is presented to overcome the difficulty of determining the shape localization in the absense of global coordinates. Thirdly, we propose the distributed shape formation control strategy utilizing only local measurements. The details are as follows.
@@ V-A1 Graphical shape representation
Similar to our previous work [8], a user-specified graphical binary image that represents the desired formation shape should be provided. As shown in Fig. 2C, the black cells correspond to the desired shape.
Each cell is described by two basic parameters $\rho$ and $\xi_{\rho}$. The coordinates $\rho=(\rho_{x},\rho_{y})$ are the column and row indexes of a cell, see Fig. 2C. The scalar $\xi_{\rho}\in[0,1]$ is the color of the cell: $\xi_{\rho}=0$ if it is black and $\xi_{\rho}=1$ if it is white. Let $n_{\mathrm{cell}}$ be the number of black cells. Therefore, the desired formation shape can be described by the index set =$\mathcal{F}=\{\rho:\xi_{\rho}=0\}$ and $\lvert\mathcal{F}\rvert=n_{\mathrm{cell}}$.
@@ V-A2 Gray transformation of the desired formation
$\xi_{\rho}^{k}=\min_{\rho^{{}^{\prime}}\in\mathcal{M}_{\rho}}\left(\xi_{\rho^{{}^{\prime}}}^{k-1}+\frac{1}{l}\right),k=1,2,...,l-1,$ (13)
where the superscript $k$ denotes the $k$-th outer layer surrounding the desired shape, as shown in Fig. 2C. The set $\mathcal{M}_{\rho}$ is composed of $3\times 3$ cells including $\rho$ and its surrounding cells. With iteration (13), the gray formation shape is $\mathcal{F}_{\mathrm{gray}}=\{\rho:\xi_{\rho}\in[0,1]\}$.
@@ V-A3 Parameters of the desired formation
The desired shape represented by image is merely graphical. Each robot should further parameterize it such that it is implementable in physical world. The size of the formation shape should be determined. The length $l_{\mathrm{cell}}$ of each cell represents the size of the formation shape and can be selected as
$l_{\mathrm{cell}}=\sqrt{\frac{\pi}{4}\frac{N}{n_{\mathrm{cell}}}}r_{\mathrm{avoid}},$
where $r_{\mathrm{avoid}}$ is the inter-robot collision avoidance distance. In this paper, we assume that the IO coordinate of each robot has the same direction such that the orientation of the desired shape can be determined as the same. Note that it is a common assumption in several relative localization results including [17] and [12], and it can be achieved through different sensors including magnetic compasses.
@@ V-B1 Neighboring robot definition
$\mathcal{N}_{i}$ $=\mathcal{N}_{i}^{\mathrm{cur}}\cap\mathcal{N}_{i}^{\mathrm{pas}},$ (14)
@@
$\mathcal{N}_{i}^{\mathrm{cur}}=\{j|\lVert p_{i}(t)-p_{j}(t)\rVert\leq r_{\mathrm{sense}}\},$ (15)
@@
$\mathcal{N}_{i}^{\mathrm{pas}}=\{j|\exists t^{\prime}\leq t,\lVert p_{i}(t^{\prime})-p_{j}(t^{\prime})\rVert\leq r_{\mathrm{neigh}}\},$ (16)
where $r_{\mathrm{sense}}$ and $r_{\mathrm{neigh}}<r_{\mathrm{sense}}$ denote the maximum measurement radius and a user-defined neighboring radius. The set $\mathcal{N}_{i}^{\mathrm{cur}}$ denotes the robots which are in the robot $i$’s sensing range at current time. The set $\mathcal{N}_{i}^{\mathrm{pas}}$ denotes the robots which are close to robot $i$ in the past and current time. The neighbor set of robot $i$ is defined as the intersection of $\mathcal{N}_{i}^{\mathrm{cur}}$ and $\mathcal{N}_{i}^{\mathrm{pas}}$. According to the definition of the neighbor set, it can be seen that when robot $j$ first become a new neighbor of robot $i$, it is within the circular range of $r_{\mathrm{neigh}}$ of robot $i$ and is unlocalized by robot $i$, i.e., robot $i$ has not collected sufficient measurement data. The purpose of defining neighbors as (14) is to maintain the measurement for a period of time between robot $i$ and $j$, which become the neighboring robots at the first time. As a result, the localization between the two robots can be achieved if the measurement data collected during this time satisfies $\mathrm{rank}(S_{ij})>0$.
As shown in Algorithm 1, each robot maintains three sets, i.e., $\mathcal{CD}_{i}$, $\mathcal{RD}_{i}$, $\mathcal{CL}_{i}$ to represent the current measurement data to its neighboring robots, the historical measurement data to neighbors and the localized robot ID. Function $\mathrm{JdgNewNeigh}(\cdot)$ is used to detect the new neighbors which are unlocalized. Function $\mathrm{JdgDt}(\cdot)$ judges whether sufficient measurement data is collected, i.e., whether $\frac{\lambda_{\min}(S_{ij})}{\lambda_{\max}(S_{ij})}$ is larger than the data matrix threshold $\lambda_{0}$. If there are new neighbors to be localized or the data collection is insufficient, the localization-enhancing velocity command $v_{i}^{\mathrm{enh}}$ will derive robot $i$ to make localization-enhancing motion within radius $r_{i}$. During this process, robot $i$ collects measurement data (see Algorithm 1, line 9 - 17). The detailed control protocol can be found in Section V-C.

### v4

@@ IV-A Relative Localization Problem
For each robot $i\in\mathcal{V}$ and $j\in\mathcal{N}_{i}$, denote the relative position between robots $i$ and $j$ at time $t_{k}=k\Delta t$ as $p_{ij}(t_{k})=p_{i}(t_{k})-p_{j}(t_{k})$, where $k\in\mathbb{N}$ is the number of samples, and $\Delta t$ is the sampling interval. Here, $p_{ij}(t_{k})$ is related to its initial relative position $p_{ij}(t_{0})$ and satisfies
$p_{ij}(t_{k})=p_{ij}(t_{0})+z_{ij}(t_{k})$ (1)
@@
According to the cosine theorem, the relative distance, relative displacement and relative position between robot $i$ and robot $j$ satisfy
$u_{ij}^{T}(t_{k})p_{ij}(t_{k})=\frac{1}{2}\left(d_{ij}^{2}(t_{k+1})-d_{ij}^{2}(t_{k})-u_{ij}^{2}(t_{k})\right)$ (2)
[…]
Define an auxiliary signal $y_{ij}(t_{k})$ as,
$y_{ij}(t_{k})$ $=u_{ij}^{T}(t_{k})p_{ij}(t_{0})$ (3)
$=u_{ij}^{T}(t_{k})p_{ij}(t_{k})-u_{ij}^{T}(t_{k})z_{ij}(t_{k})$
$=\frac{1}{2}\left(d_{ij}^{2}(t_{k+1})-d_{ij}^{2}(t_{k})-u_{ij}^{2}(t_{k})\right)-u_{ij}^{T}(t_{k})z_{ij}(t_{k}).$
One can see that $y_{ij}(t_{k})$ is available, since both $u_{ij}$ and $z_{ij}$ are directly obtainable. With the fact that both $y_{ij}(t_{k})$ and $u_{ij}(t_{k})$ can be measured, the adaptive parameter update law can be designed to estimate the initial relative position $p_{ij}(t_{0})$. As a result, the relative localization task can be achieved.
Denote $\hat{p}_{ij,0}(t_{k})$ as an estimation of the initial relative position $p_{ij}(t_{0})$ and $\tilde{p}_{ij,0}(t_{k})=\hat{p}_{ij,0}(t_{k})-p_{ij}(t_{0})$ as the estimation error. We can define the innovation $\epsilon_{ij}(t_{k})$ as
@@
The innovation $\epsilon_{ij}(t_{k})$ will be used later to design the concurrent-learning based estimator in Section IV-C. Based on the estimation problem modeled as (3), methods proposed in [8], [13] and [37] attempt to estimate the initial relative position $p_{ij}(t_{0})$. However, all these existing results require that relative displacement $u_{ij}(t)$ satisfies the PE condition. To satisfy this condition, the persistent noise or periodic signal is adopted in the control command of the robot, which is not conducive to the expansion of massive robot swarms.
@@ IV-B Data Collection Strategy
Figure 3: Geometric relationship between the displacements and distance measurements of the two robots. A. One sampling interval $\Delta t$ case. B. A sketch for $h$ sampling interval $\Delta t$ case
In the next section, how to collect varables $d_{ij}(t_{k})$, $d_{ij}(t_{k+1})$ and $u_{ij}(t_{k})$ to determine $p_{ij}(t_{0})$ will be presented. However, for the real IO sensor, if the sampling interval $\Delta t$ is too small, the measurements of $u_{ij}(t_{k})$ may be drowned in noise in practice. To address this issue, we propose the following data collection strategy.
Whenever two robots start localization, a data collection takes place between them. As shown in Fig. 3B, assuming that a data collection starts at time instant $t_{k}$, then it will end at $t_{k+h}=t_{k}+h\Delta t$. For biref, we denote $t_{{\rm c},m}$ and $t_{{\rm c},m+1}$ as the starting and ending time of the $m$-th ($m=1,\ldots,\varsigma$) data collection. Note that $\varsigma$ is the data collection number and $t_{{\rm c},m+1}$ is also the starting time of the $(m+1)$-th data collection. The interval between the two consecutive collections is $h\Delta t,h\in\mathbb{N}$, i.e., $t_{{\rm c},m+1}-t_{{\rm c},m}=h\Delta t$. The whole data collection process begins at $t_{{\rm c},1}$ and terminates at $t_{{\rm c},\varsigma}$. The heuristic values of $h$ and $\varsigma$ will be discussed later in Section IV-D. In this way, the innovation $\epsilon_{ij}(t_{k})$ in (4) can be rewritten as
$\epsilon_{ij}(t_{{\rm c},m})=$ $u_{ij}^{T}(t_{{\rm c},m})\hat{p}_{ij,0}(t_{k})-y_{ij}(t_{{\rm c},m})$ (5)
$y_{ij}(t_{{\rm c},m})=$ $\frac{1}{2}\left(d_{ij}^{2}(t_{c,m+1})-d_{ij}^{2}(t_{{\rm c},m})-u_{ij}^{2}(t_{{\rm c},m})\right)$
$-u_{ij}^{T}(t_{{\rm c},m})z_{ij}(t_{{\rm c},m})$
where $z_{ij}(t_{{\rm c},m})=\sum_{\iota=1}^{m}u_{ij}(t_{{\rm c},\iota})$. The innovation $\epsilon_{ij}(t_{{\rm c},m})$ will be used later in Section IV-C. Note that the tandem data collection strategy is helpful for increasing the modulus of the relative displacement $u_{ij}(t_{{\rm c},m})$, which is beneficial for reducing relative error in noisy measurements. This trick helps to improve estimation accuracy in real applications.
@@ IV-C Concurrent-Learning based Estimator
The concurrent-learning based estimator $\hat{p}_{ij,0}(t_{k})$, which is used to estimate $p_{ij}(t_{0})$, can be designed as
@@
$\eta=\frac{\lambda_{\min}(S_{ij})}{(\lambda_{\max}(U_{ij}(t_{k}))+\lambda_{\max}(S_{ij}))^{2}}$ (7)
where matrices $U_{ij}(t_{k})$ and $S_{ij}$ are defined as $U_{ij}(t_{k})=u_{ij}(t_{k})u_{ij}^{T}(t_{k})\in\mathbb{R}^{2\times 2}$ and $S_{ij}=R_{ij}R_{ij}^{T}\in\mathbb{R}^{2\times 2}$, respectively. Here, $R_{ij}=[u_{ij}(t_{c,1}),...,u_{ij}(t_{c,\varsigma})]$. Note that (6) represents the updates at time instant $t_{k+1}$ and employs both the present measurements $d_{ij}(t_{k+1})$, $d_{ij}(t_{k})$, $u_{ij}(t_{k})$ and the historical measurements $d_{ij}(t_{{\rm c},m})$, $d_{ij}(t_{c,m+1})$, $u_{ij}(t_{{\rm c},m})$ in its calculations. All these measurements are available at $t_{k+1}$.
@@
Recall that $z_{ij}(t)$ is the relative displacement which is available through the IO measurement and inter-robot communication. For actual sensors, UWB measurement often has low frame rates, while the frequency of IO is often high. So we model the measurement system as a discrete system. On the other hand, when we estimate the real-time relative position of the robot based on Eq. (1), we can use the IO measurement which is approximately continuous. In the next theorem, we will show that the convergence of the proposed estimator (6) can be guaranteed if the recorded data matrix $S_{ij}$ satisfies the following assumption.
@@ IV-D Localization Enhancement Motion
In real applications, nevertheless, we still need to address the questions on how to design the relative motion trajectories of two robots and how to collect the measurement data, i.e. how to choose $\varsigma$ and $h$ such that the convergence speed of estimator (6) is the largest. According to Theorem 1, to improve the convergence speed of the estimator, the ratio of the minimum and maximum eigenvalue $\frac{\lambda_{\min}(S_{ij})}{\lambda_{\max}(S_{ij})}$ of the recorded data matrix $S_{ij}$ should be as large as desired. To achieve this, we design the velocity control command of each robot $i$ as the following circular motion velocity
$v_{i}$ $=[r_{i}\mathrm{cos}(w_{i}t),r_{i}\mathrm{sin}(w_{i}t)]$ (10)
where $w_{i}=1/i$ is the designed angular velocity and $r_{i}$ is the motion radius. The reason for this design is to enable the estimator (6) to have a faster convergence speed. For robots $i$ and $j$, we select the measurement interval time $h\Delta t$ and the data set number $\varsigma$ which satisfy
$\mathrm{mod}(2\pi,h\Delta t)$ $=0$ (11a)
$\varsigma$ $=ij\varsigma_{0}$ (11b)
@@
Theorem 2 shows that with careful design of the relative motion of the two robots, the proposed estimator has the largest convergence rate. In the practical applications, the ratio $\frac{\lambda_{\min}(S_{ij})}{\lambda_{\max}(S_{ij})}$ can be calculated at each data measurement time instant $t_{{\rm c},m}$. Define the threshold $\lambda_{0}$ of $\frac{\lambda_{\min}(S_{ij})}{\lambda_{\max}(S_{ij})}$. When $\frac{\lambda_{\min}(S_{ij})}{\lambda_{\max}(S_{ij})}$ is greater than $\lambda_{0}$, the data collection is sufficient, there is no need to maintain the relative motion between the two robots and the estimator will converge exponentially. This result directly supports our design of localization enhance command in Section VI B.
@@ V Behavior-Based Shape Formation
Previously, we have shown how to achieve relative localization with only distance and displacement measurements. In this section, we will address the problem of how to achieve shape formation control through behavior based control strategy. To achieve this, we first define and parameterize the desired formation shape. Then, a consensus based shape localization agreement protocol is presented to determine the shape localization in the absence of global coordinates. Thirdly, we propose the distributed shape formation control strategy utilizing local measurements. The details are as follows.
@@ V-A1 Graphical shape representation
Similar to our previous work [5], a user-specified graphical binary image should be provided to represent the desired formation shape. As shown in Fig. 2C, the black cells correspond to the desired shape. Each cell is described by two basic parameters $\rho$ and $\xi_{\rho}$. The coordinates $\rho=(\rho_{x},\rho_{y})$ are the column and row indexes of a cell, see Fig. 2C. The scalar $\xi_{\rho}\in[0,1]$ is the color of the cell: $\xi_{\rho}=0$ if it is black and $\xi_{\rho}=1$ if it is white. Let $n_{\mathrm{cell}}$ denote the black cell number in the binary image. Generally, the larger the value of $n_{\mathrm{cell}}$, the clearer the target image, but it also consumes more storage space. Therefore, the desired formation shape can be described by the index set =$\mathcal{F}=\{\rho:\xi_{\rho}=0\}$ and $\lvert\mathcal{F}\rvert=n_{\mathrm{cell}}$.
@@ V-A2 Gray transformation of the desired formation
$\xi_{\rho}^{k}=\min_{\rho^{{}^{\prime}}\in\mathcal{M}_{\rho}}\left(\xi_{\rho^{{}^{\prime}}}^{k-1}+\frac{1}{l}\right),k=1,2,...,l-1$ (12)
where the superscript $k$ denotes the $k$-th outer layer surrounding the grid, as shown in Fig. 2C. The set $\mathcal{M}_{\rho}$ is composed of $3\times 3$ cells including $\rho$ and its surrounding cells. With iteration (12), the gray formation shape is $\mathcal{F}_{\mathrm{gray}}=\{\rho:\xi_{\rho}\in[0,1]\}$.
@@ V-A3 Parameters of the desired formation
The desired shape represented by image is merely graphical. Each robot should further parameterize it such that it is implementable in physical world. The size of the formation shape should be determined. The length $l_{\mathrm{cell}}$ of each cell represents the size of the formation shape and can be selected as $l_{\mathrm{cell}}=\sqrt{\frac{\pi}{4}\frac{N}{n_{\mathrm{cell}}}}r_{\mathrm{avoid}}$, where $r_{\mathrm{avoid}}$ is the inter-robot collision avoidance distance. In this article, we assume that the IO coordinate of each robot has the same direction such that the orientation of the desired shape can be determined as the same. Note that it is a common assumption in several relative localization results including [13] and [8], and it can be achieved through different sensors including magnetic compasses.
@@ V-B1 Neighboring robot definition
$\mathcal{N}_{i}$ $=\mathcal{N}_{i}^{\mathrm{cur}}\cap\mathcal{N}_{i}^{\mathrm{pas}}$ (13)
@@
$\mathcal{N}_{i}^{\mathrm{cur}}$ =$=\{j|\lVert p_{i}(t)-p_{j}(t)\rVert\leq r_{\mathrm{sense}},j\neq i\}$
@@
$\mathcal{N}_{i}^{\mathrm{pas}}$ =$=\{j|\exists t^{\prime}\leq t,\lVert p_{i}(t^{\prime})-p_{j}(t^{\prime})\rVert\leq r_{\mathrm{neigh}},j\neq i\}$
where $r_{\mathrm{sense}}$ and $r_{\mathrm{neigh}}<r_{\mathrm{sense}}$ denote the maximum measurement radius and a user-defined neighboring radius. The set $\mathcal{N}_{i}^{\mathrm{cur}}$ denotes the robots which are in the robot $i$’s sensing range at the current time. The set $\mathcal{N}_{i}^{\mathrm{pas}}$ denotes the robots which are close to robot $i$ in the past and current time. The neighbor set of robot $i$ is defined as the intersection of $\mathcal{N}_{i}^{\mathrm{cur}}$ and $\mathcal{N}_{i}^{\mathrm{pas}}$. According to the definition of the neighbor set, it can be seen that when robot $j$ firstly becomes a new neighbor of robot $i$, it is within the circular range of $r_{\mathrm{neigh}}$ of robot $i$ and is unlocalized by robot $i$, i.e., robot $i$ has not collected sufficient measurement data. Then they need to collect measurement data to achieve relative localization. However, if we directly use whether the distance between robots is less than $r_{\mathrm{sense}}$ as the basis for judging neighbors, the distance $d_{ij}$ between robots $i$ and $j$ may exceed $r_{\mathrm{sense}}$ when robots collect data on relative motion. To maintain the measurement for a period of time between robots $i$ and $j$ when they firstly meet each other, we define the set of neighbors as (13). One can see that communication and measurement between robots $i$ and $j$ can be maintained at least for a time duartion $\frac{r_{\mathrm{sense}}-r_{\mathrm{neigh}}}{2v_{\mathrm{max}}}$.
As shown in Algorithm 1, each robot maintains three sets, i.e., $\mathcal{CD}_{i}$, $\mathcal{RD}_{i}$, $\mathcal{CL}_{i}$ to represent the current measurement data to its neighboring robots, the historical measurement data to neighbors and the localized robot ID. Function $\mathrm{JdgNewNeigh}(\cdot)$ is used to detect the new neighbors which are unlocalized. Function $\mathrm{JdgDt}(\cdot)$ judges whether sufficient measurement data is collected, i.e., whether $\frac{\lambda_{\min}(S_{ij})}{\lambda_{\max}(S_{ij})}$ is larger than the data matrix threshold $\lambda_{0}$. If there are new neighbors to be localized or the data collection is insufficient, the localization-enhancing velocity command $v_{i}^{\mathrm{enh}}$ will derive robot $i$ to make localization-enhancing motion within radius $r_{i}$. During this process, robot $i$ collects measurement data (see Algorithm 1, lines 9 - 17). The detailed control protocol can be found in Section V-C.